# Can Transformers Learn Full Bayesian Inference in Context?

### v1

@@ Appendix I Ablation: Different learning rates for VI
Laplace Approximation 1e-2 1.000 ($\pm$ 0.000) 4.380 (±($\pm$( ± 1.386) 4.838 ($\pm$ 1.521) 1.000 ($\pm$ 0.000) 4.588 (±($\pm$( ± 1.229) 6.813 ($\pm$ 1.697)
Laplace Approximation 1e-3 1.000 ($\pm$ 0.000) 3.893 (±($\pm$( ± 1.433) 4.010 ($\pm$ 1.233) 1.000 ($\pm$ 0.000) 4.699 (±($\pm$( ± 1.193) 6.986 ($\pm$ 0.981)
Laplace Approximation 1e-4 1.000 ($\pm$ 0.000) 4.463 (±($\pm$( ± 1.117) 4.610 ($\pm$ 1.027) 1.000 ($\pm$ 0.000) 4.710 (±($\pm$( ± 1.205) 6.995 ($\pm$ 0.869)
VI: DiagonalNormal 1e-2 0.979 ($\pm$ 0.138) 1.370 (±($\pm$( ± 1.394) 3.522 ($\pm$ 1.634) 0.985 ($\pm$ 0.030) 2.384 (±($\pm$( ± 1.318) 6.202 ($\pm$ 1.747)
VI: DiagonalNormal 1e-3 0.990 ($\pm$ 0.096) 1.454 (±($\pm$( ± 1.454) 3.650 ($\pm$ 1.743) 0.999 ($\pm$ 0.002) 3.026 (±($\pm$( ± 0.977) 6.959 ($\pm$ 0.890)
VI: DiagonalNormal 1e-4 1.000 ($\pm$ 0.001) 2.390 (±($\pm$( ± 1.177) 4.903 ($\pm$ 1.278) 0.998 ($\pm$ 0.007) 2.830 (±($\pm$( ± 1.001) 7.007 ($\pm$ 0.987)
VI: MultivariateNormal 1e-2 0.978 ($\pm$ 0.119) 1.351 (±($\pm$( ± 1.410) 3.474 ($\pm$ 1.604) 0.987 ($\pm$ 0.024) 2.375 (±($\pm$( ± 1.304) 6.189 ($\pm$ 1.761)
VI: MultivariateNormal 1e-3 0.980 ($\pm$ 0.089) 1.476 (±($\pm$( ± 1.480) 3.681 ($\pm$ 1.734) 0.997 ($\pm$ 0.008) 2.808 (±($\pm$( ± 1.014) 6.964 ($\pm$ 0.944)
VI: MultivariateNormal 1e-4 1.000 ($\pm$ 0.001) 2.114 (±($\pm$( ± 1.140) 4.532 ($\pm$ 1.187) 0.997 ($\pm$ 0.007) 2.799 (±($\pm$( ± 1.012) 6.963 ($\pm$ 0.950)
VI: Structured Normal 1e-2 0.958 ($\pm$ 0.129) 1.246 (±($\pm$( ± 1.615) 3.225 ($\pm$ 1.701) 1.000 ($\pm$ 0.001) 2.911 (±($\pm$( ± 0.753) 6.675 ($\pm$ 1.403)
VI: Structured Normal 1e-3 0.979 ($\pm$ 0.092) 1.593 (±($\pm$( ± 1.561) 3.395 ($\pm$ 1.440) 0.998 ($\pm$ 0.007) 2.882 (±($\pm$( ± 1.070) 6.968 ($\pm$ 0.941)
VI: Structured Normal 1e-4 1.000 ($\pm$ 0.001) 2.270 (±($\pm$( ± 1.133) 4.733 ($\pm$ 1.162) 0.997 ($\pm$ 0.009) 2.802 (±($\pm$( ± 1.012) 6.953 ($\pm$ 0.948)
VI: IAF 1e-2 0.998 ($\pm$ 0.003) 1.539 (±($\pm$( ± 0.691) 8.371 ($\pm$ 0.750) 0.987 ($\pm$ 0.022) 1.376 (±($\pm$( ± 0.799) 8.082 ($\pm$ 1.352)
VI: IAF 1e-3 0.997 ($\pm$ 0.004) 1.443 (±($\pm$( ± 0.564) 8.517 ($\pm$ 0.820) 0.988 ($\pm$ 0.020) 1.304 (±($\pm$( ± 0.855) 8.425 ($\pm$ 1.281)
VI: IAF 1e-4 0.997 ($\pm$ 0.004) 1.602 (±($\pm$( ± 0.628) 7.888 ($\pm$ 0.783) 0.987 ($\pm$ 0.020) 1.380 (±($\pm$( ± 0.848) 7.729 ($\pm$ 1.322)
Appendix J Preprocessing of the real-world datasets
Appendix K Background on conditional flow-matching

### v2

@@ I.3 Gaussian Mixture Models
ICL (ours) 0.760 ($\pm$ 0.092) 0.303 ($\pm$ 0.548) 2.095 ($\pm$ 1.692) 0.847 ($\pm$ 0.082) 0.486 ($\pm$ 0.623) 4.054 ($\pm$ 2.782)
ICL (ours) 0.812 ($\pm$ 0.061) 0.159 ($\pm$ 0.154) 2.314 ($\pm$ 0.926) 0.937 ($\pm$ 0.041) 0.282 ($\pm$ 0.131) 3.947 ($\pm$ 1.055)
ICL (ours) 1.000 ($\pm$ 0.000) 0.582 ($\pm$ 0.280) 8.708 ($\pm$ 4.945) 1.000 ($\pm$ 0.000) 1.869 ($\pm$ 0.342) 33.230 ($\pm$ 8.095)
Scenario 4 SGLD 1.000 ($\pm$ 0.000) 3.464 $(\pm$ 1.098) 6.995 ($\pm$ 5.554) 1.000 ($\pm$ 0.000) 2.555 $(\pm$ 0.494) 9.477 ($\pm$ 3.432)
ICL (ours) 1.000 ($\pm$ 0.000) 2.451 ($\pm$ 0.868) 8.333 ($\pm$ 4.202) 1.000 ($\pm$ 0.000) 2.518 ($\pm$ 0.694) 11.938 ($\pm$ 2.956)
For GMMs (Table 44), ICL demonstrates a clear advantage in most scenarios. In Scenario 1, ICL reduces synthetic C2ST from 1.000 to 0.760 and real-world $\mathcal{W}_{2}$ from 6.510 to 4.054. Scenario 2 shows synthetic C2ST improving from 1.000 to 0.812, and MMD from 3.046 to 0.159. While in scenarios 3, ICL has a singificantly lower MMD score on the synthetic data, the other differences are not significant.
Appendix Q Evaluation the choice of classifier for the C2ST metric
In this section, we validate the choice of the classifier for the C2ST metric by comparing the ROC characteristic of a random forest (our default choice) and a neural network in distinguishing posterior samples. In summary, we find that despite minor differences, the two metrics yield the same overall results. Across all scenarios, both Random Forest (RF) and Neural Network NN classifiers yield quite consistent rankings of model performance with only insubstantial deviations in terms of the big picture. In particular, ICL is consistently among the top-performing approaches under both evaluation metrics. Out of the 14 total scenario–domain combinations (7 scenarios × 2 dataset types), the RF and NN metrics identify the same best-performing model in 12 cases.
Table 45: Generalized Linear Models: Comparison of C2ST scores with a Random Forest (RF) and a Neural Network (NN). For the NN we follow the setup of Lueckmann et al., 2021. Evaluation across seven distinct scenarios on 50 synthetic and 17 real-world datasets. All results within two standard errors of the best average result in each scenario are marked in bold.
Scenario Model Synthetic Evaluation Real-World Evaluation
C2ST RF ($\downarrow$) C2ST NN ($\downarrow$) C2ST RF ($\downarrow$) C2ST NN ($\downarrow$)
Scenario 1 Laplace Approximation 1.000 ($\pm$ 0.000) 0.998 ($\pm$ 0.000) 1.000 ($\pm$ 0.000) 0.998 ($\pm$ 0.000)
VI: DiagonalNormal 0.904 ($\pm$ 0.076) 0.857 ($\pm$ 0.001) 0.797 ($\pm$ 0.083) 0.803 ($\pm$ 0.004)
VI: MultivariateNormal 0.750 ($\pm$ 0.128) 0.780 ($\pm$ 0.002) 0.607 ($\pm$ 0.070) 0.713 ($\pm$ 0.004)
VI: Structured Normal 0.753 ($\pm$ 0.126) 0.781 ($\pm$ 0.002) 0.600 ($\pm$ 0.070) 0.705 ($\pm$ 0.004)
VI: IAF 0.777 ($\pm$ 0.122) 0.793 ($\pm$ 0.002) 0.683 ($\pm$ 0.132) 0.746 ($\pm$ 0.006)
HMC 0.745 ($\pm$ 0.130) 0.777 ($\pm$ 0.002) 0.595 ($\pm$ 0.075) 0.702 ($\pm$ 0.004)
ICL (ours) 0.765 ($\pm$ 0.123) 0.712 ($\pm$ 0.002) 0.614 ($\pm$ 0.074) 0.701 ($\pm$ 0.004)
Scenario 2 Laplace Approximation 1.000 ($\pm$ 0.000) 0.998 ($\pm$ 0.000) 1.000 ($\pm$ 0.000) 0.998 ($\pm$ 0.000)
VI: DiagonalNormal 0.957 ($\pm$ 0.091) 0.883 ($\pm$ 0.002) 0.892 ($\pm$ 0.044) 0.851 ($\pm$ 0.003)
VI: MultivariateNormal 0.910 ($\pm$ 0.131) 0.860 ($\pm$ 0.002) 0.820 ($\pm$ 0.031) 0.815 ($\pm$ 0.003)
VI: Structured Normal 0.908 ($\pm$ 0.119) 0.859 ($\pm$ 0.002) 0.824 ($\pm$ 0.023) 0.817 ($\pm$ 0.003)
VI: IAF 0.968 ($\pm$ 0.063) 0.889 ($\pm$ 0.001) 0.888 ($\pm$ 0.067) 0.849 ($\pm$ 0.004)
ICL (ours) 0.839 ($\pm$ 0.072) 0.824 ($\pm$ 0.001) 0.768 ($\pm$ 0.033) 0.789 ($\pm$ 0.003)
Scenario 3 Laplace Approximation 1.000 ($\pm$ 0.000) 0.998 ($\pm$ 0.000) 1.000 ($\pm$ 0.000) 0.998 ($\pm$ 0.000)
VI: DiagonalNormal 0.866 ($\pm$ 0.101) 0.838 ($\pm$ 0.002) 0.797 ($\pm$ 0.083) 0.803 ($\pm$ 0.004)
VI: MultivariateNormal 0.656 ($\pm$ 0.131) 0.733 ($\pm$ 0.002) 0.590 ($\pm$ 0.035) 0.685 ($\pm$ 0.003)
VI: Structured Normal 0.653 ($\pm$ 0.125) 0.731 ($\pm$ 0.002) 0.582 ($\pm$ 0.028) 0.681 ($\pm$ 0.003)
VI: IAF 0.751 ($\pm$ 0.148) 0.780 ($\pm$ 0.002) 0.673 ($\pm$ 0.141) 0.741 ($\pm$ 0.006)
ICL (ours) 0.611 ($\pm$ 0.070) 0.710 ($\pm$ 0.001) 0.576 ($\pm$ 0.027) 0.693 ($\pm$ 0.003)
Scenario 4 Laplace Approximation 1.000 ($\pm$ 0.000) 0.998 ($\pm$ 0.000) 1.000 ($\pm$ 0.000) 0.998 ($\pm$ 0.000)
VI: DiagonalNormal 0.968 ($\pm$ 0.036) 0.889 ($\pm$ 0.001) 0.916 ($\pm$ 0.040) 0.863 ($\pm$ 0.003)
VI: MultivariateNormal 0.855 ($\pm$ 0.123) 0.832 ($\pm$ 0.002) 0.771 ($\pm$ 0.017) 0.790 ($\pm$ 0.002)
VI: Structured Normal 0.847 ($\pm$ 0.116) 0.828 ($\pm$ 0.002) 0.769 ($\pm$ 0.012) 0.789 ($\pm$ 0.002)
VI: IAF 0.942 ($\pm$ 0.077) 0.876 ($\pm$ 0.001) 0.833 ($\pm$ 0.069) 0.821 ($\pm$ 0.004)
ICL (ours) 0.753 ($\pm$ 0.049) 0.781 ($\pm$ 0.001) 0.762 ($\pm$ 0.015) 0.786 ($\pm$ 0.002)
Scenario 5 Laplace Approximation 1.000 ($\pm$ 0.000) 0.998 ($\pm$ 0.000) 1.000 ($\pm$ 0.000) 0.998 ($\pm$ 0.000)
VI: DiagonalNormal 0.866 ($\pm$ 0.085) 0.838 ($\pm$ 0.002) 0.810 ($\pm$ 0.036) 0.810 ($\pm$ 0.003)
VI: MultivariateNormal 0.765 ($\pm$ 0.100) 0.787 ($\pm$ 0.002) 0.711 ($\pm$ 0.038) 0.760 ($\pm$ 0.003)
VI: Structured Normal 0.758 ($\pm$ 0.098) 0.784 ($\pm$ 0.002) 0.705 ($\pm$ 0.032) 0.757 ($\pm$ 0.003)
VI: IAF 0.814 ($\pm$ 0.105) 0.812 ($\pm$ 0.002) 0.777 ($\pm$ 0.106) 0.793 ($\pm$ 0.005)
ICL (ours) 0.621 ($\pm$ 0.063) 0.715 ($\pm$ 0.001) 0.610 ($\pm$ 0.045) 0.710 ($\pm$ 0.003)
Scenario 6 Laplace Approximation 1.000 ($\pm$ 0.000) 0.998 ($\pm$ 0.000) 1.000 ($\pm$ 0.000) 0.998 ($\pm$ 0.000)
VI: DiagonalNormal 0.724 ($\pm$ 0.060) 0.767 ($\pm$ 0.001) 0.703 ($\pm$ 0.039) 0.756 ($\pm$ 0.003)
VI: MultivariateNormal 0.534 ($\pm$ 0.018) 0.672 ($\pm$ 0.001) 0.538 ($\pm$ 0.019) 0.674 ($\pm$ 0.002)
VI: Structured Normal 0.536 ($\pm$ 0.016) 0.673 ($\pm$ 0.001) 0.536 ($\pm$ 0.019) 0.673 ($\pm$ 0.002)
VI: IAF 0.542 ($\pm$ 0.026) 0.676 ($\pm$ 0.001) 0.535 ($\pm$ 0.015) 0.672 ($\pm$ 0.002)
ICL (ours) 0.532 ($\pm$ 0.019) 0.671 ($\pm$ 0.001) 0.556 ($\pm$ 0.017) 0.653 ($\pm$ 0.002)
Scenario 7 Laplace Approximation 1.000 ($\pm$ 0.000) 0.998 ($\pm$ 0.000) 1.000 ($\pm$ 0.000) 0.998 ($\pm$ 0.000)
VI: DiagonalNormal 0.938 ($\pm$ 0.074) 0.874 ($\pm$ 0.001) 0.936 ($\pm$ 0.024) 0.873 ($\pm$ 0.003)
VI: MultivariateNormal 0.814 ($\pm$ 0.181) 0.812 ($\pm$ 0.002) 0.741 ($\pm$ 0.020) 0.775 ($\pm$ 0.003)
VI: Structured Normal 0.824 ($\pm$ 0.177) 0.817 ($\pm$ 0.002) 0.734 ($\pm$ 0.025) 0.772 ($\pm$ 0.003)
VI: IAF 0.939 ($\pm$ 0.091) 0.874 ($\pm$ 0.002) 0.864 ($\pm$ 0.093) 0.837 ($\pm$ 0.005)
ICL (ours) 0.700 ($\pm$ 0.116) 0.721 ($\pm$ 0.002) 0.773 ($\pm$ 0.048) 0.751 ($\pm$ 0.003)
Table 46: Factor Analysis: Comparison of C2ST scores using a Random Forest (RF) and a Neural Network (NN) classifier across six different scenarios on 50 synthetic and 17 real-world datasets. For the NN we follow the setup of Lueckmann et al., 2021. All results within two standard errors of the best average result in each scenario are marked in bold.
Scenario Model Synthetic Evaluation Real-World Evaluation
C2ST RF ($\downarrow$) C2ST NN ($\downarrow$) C2ST RF ($\downarrow$) C2ST NN ($\downarrow$)
Scenario 1 Laplace Approximation 1.000 ($\pm$ 0.000) 0.997 ($\pm$ 0.000) 1.000 ($\pm$ 0.000) 0.997 ($\pm$ 0.000)
VI: DiagonalNormal 1.000 ($\pm$ 0.001) 0.997 ($\pm$ 0.000) 0.979 ($\pm$ 0.008) 0.950 ($\pm$ 0.001)
VI: MultivariateNormal 0.998 ($\pm$ 0.003) 0.960 ($\pm$ 0.000) 0.966 ($\pm$ 0.010) 0.944 ($\pm$ 0.001)
VI: Structured Normal 0.997 ($\pm$ 0.004) 0.959 ($\pm$ 0.000) 0.979 ($\pm$ 0.010) 0.950 ($\pm$ 0.001)
VI: IAF 0.953 ($\pm$ 0.104) 0.937 ($\pm$ 0.001) 0.849 ($\pm$ 0.075) 0.885 ($\pm$ 0.003)
ICL (ours) 0.552 ($\pm$ 0.028) 0.737 ($\pm$ 0.000) 0.606 ($\pm$ 0.038) 0.764 ($\pm$ 0.001)
Scenario 2 Laplace Approximation 1.000 ($\pm$ 0.000) 0.997 ($\pm$ 0.000) 1.000 ($\pm$ 0.000) 0.997 ($\pm$ 0.000)
VI: DiagonalNormal 0.998 ($\pm$ 0.002) 0.960 ($\pm$ 0.000) 0.975 ($\pm$ 0.010) 0.948 ($\pm$ 0.001)
VI: MultivariateNormal 0.989 ($\pm$ 0.009) 0.955 ($\pm$ 0.000) 0.951 ($\pm$ 0.025) 0.936 ($\pm$ 0.001)
VI: Structured Normal 0.984 ($\pm$ 0.031) 0.953 ($\pm$ 0.000) 0.958 ($\pm$ 0.025) 0.940 ($\pm$ 0.001)
VI: IAF 0.966 ($\pm$ 0.066) 0.944 ($\pm$ 0.001) 0.799 ($\pm$ 0.058) 0.860 ($\pm$ 0.002)
ICL (ours) 0.542 ($\pm$ 0.006) 0.732 ($\pm$ 0.000) 0.622 ($\pm$ 0.032) 0.772 ($\pm$ 0.001)
Scenario 3 Laplace Approximation 1.000 ($\pm$ 0.000) 0.997 ($\pm$ 0.000) 1.000 ($\pm$ 0.000) 0.997 ($\pm$ 0.000)
VI: DiagonalNormal 0.999 ($\pm$ 0.002) 0.960 ($\pm$ 0.000) 0.951 ($\pm$ 0.007) 0.936 ($\pm$ 0.001)
VI: MultivariateNormal 0.994 ($\pm$ 0.007) 0.958 ($\pm$ 0.000) 0.945 ($\pm$ 0.007) 0.933 ($\pm$ 0.001)
VI: Structured Normal 0.997 ($\pm$ 0.003) 0.959 ($\pm$ 0.000) 0.942 ($\pm$ 0.009) 0.932 ($\pm$ 0.001)
VI: IAF 0.990 ($\pm$ 0.011) 0.987 ($\pm$ 0.000) 0.928 ($\pm$ 0.015) 0.925 ($\pm$ 0.001)
ICL (ours) 0.537 ($\pm$ 0.023) 0.729 ($\pm$ 0.000) 0.609 ($\pm$ 0.019) 0.765 ($\pm$ 0.001)
Scenario 4 Laplace Approximation 1.000 ($\pm$ 0.000) 0.997 ($\pm$ 0.000) 1.000 ($\pm$ 0.000) 0.997 ($\pm$ 0.000)
VI: DiagonalNormal 1.000 ($\pm$ 0.000) 0.997 ($\pm$ 0.000) 0.977 ($\pm$ 0.003) 0.949 ($\pm$ 0.000)
VI: MultivariateNormal 0.999 ($\pm$ 0.001) 0.960 ($\pm$ 0.000) 0.973 ($\pm$ 0.008) 0.947 ($\pm$ 0.001)
VI: Structured Normal 1.000 ($\pm$ 0.000) 0.997 ($\pm$ 0.000) 0.973 ($\pm$ 0.007) 0.947 ($\pm$ 0.001)
VI: IAF 0.999 ($\pm$ 0.001) 0.960 ($\pm$ 0.000) 0.961 ($\pm$ 0.018) 0.941 ($\pm$ 0.001)
ICL (ours) 0.684 ($\pm$ 0.060) 0.803 ($\pm$ 0.001) 0.988 ($\pm$ 0.003) 0.955 ($\pm$ 0.000)
Scenario 5 Laplace Approximation 1.000 ($\pm$ 0.000) 0.997 ($\pm$ 0.000) 1.000 ($\pm$ 0.000) 0.997 ($\pm$ 0.000)
VI: DiagonalNormal 0.999 ($\pm$ 0.002) 0.960 ($\pm$ 0.000) 0.944 ($\pm$ 0.010) 0.933 ($\pm$ 0.001)
VI: MultivariateNormal 0.995 ($\pm$ 0.007) 0.958 ($\pm$ 0.000) 0.930 ($\pm$ 0.017) 0.926 ($\pm$ 0.001)
VI: Structured Normal 0.998 ($\pm$ 0.005) 0.960 ($\pm$ 0.000) 0.934 ($\pm$ 0.011) 0.928 ($\pm$ 0.001)
VI: IAF 0.992 ($\pm$ 0.012) 0.957 ($\pm$ 0.000) 0.910 ($\pm$ 0.011) 0.916 ($\pm$ 0.001)
ICL (ours) 0.535 ($\pm$ 0.016) 0.728 ($\pm$ 0.000) 0.886 ($\pm$ 0.017) 0.904 ($\pm$ 0.001)
Scenario 6 Laplace Approximation 1.000 ($\pm$ 0.000) 0.997 ($\pm$ 0.000) 1.000 ($\pm$ 0.000) 0.997 ($\pm$ 0.000)
VI: DiagonalNormal 0.998 ($\pm$ 0.002) 0.960 ($\pm$ 0.000) 0.949 ($\pm$ 0.008) 0.935 ($\pm$ 0.001)
VI: MultivariateNormal 0.991 ($\pm$ 0.013) 0.956 ($\pm$ 0.000) 0.938 ($\pm$ 0.009) 0.930 ($\pm$ 0.001)
VI: Structured Normal 0.997 ($\pm$ 0.005) 0.959 ($\pm$ 0.000) 0.944 ($\pm$ 0.006) 0.933 ($\pm$ 0.001)
VI: IAF 0.989 ($\pm$ 0.029) 0.955 ($\pm$ 0.000) 0.865 ($\pm$ 0.027) 0.893 ($\pm$ 0.001)
ICL (ours) 0.543 ($\pm$ 0.021) 0.732 ($\pm$ 0.000) 0.666 ($\pm$ 0.020) 0.794 ($\pm$ 0.001)
Table 47: Gaussian Mixture Models: Comparison of C2ST scores using a Random Forest (RF) and a Neural Network (NN) classifier across six distinct scenarios on 50 synthetic and 17 real-world datasets. All results within two standard errors of the best average result in each scenario are marked in bold. For the NN we follow the setup of Lueckmann et al., 2021. Both RF and NN classifiers yield consistent rankings, with ICL emerging as the top method in scenarios with more pronounced model mismatch.
Scenario Model Synthetic Evaluation Real-World Evaluation
C2ST RF ($\downarrow$) C2ST NN ($\downarrow$) C2ST RF ($\downarrow$) C2ST NN ($\downarrow$)
Scenario 1 Laplace Approximation 1.000 ($\pm$ 0.000) 0.997 ($\pm$ 0.000) 1.000 ($\pm$ 0.000) 0.997 ($\pm$ 0.000)
VI: DiagonalNormal 0.988 ($\pm$ 0.013) 1.012 ($\pm$ 0.000) 0.995 ($\pm$ 0.006) 0.996 ($\pm$ 0.001)
VI: MultivariateNormal 0.988 ($\pm$ 0.013) 1.012 ($\pm$ 0.000) 0.994 ($\pm$ 0.007) 0.993 ($\pm$ 0.001)
VI: Structured Normal 0.987 ($\pm$ 0.015) 0.982 ($\pm$ 0.000) 0.993 ($\pm$ 0.009) 0.992 ($\pm$ 0.001)
VI: IAF 0.989 ($\pm$ 0.013) 0.983 ($\pm$ 0.000) 0.995 ($\pm$ 0.010) 0.996 ($\pm$ 0.001)
ICL (ours) 0.760 ($\pm$ 0.092) 0.825 ($\pm$ 0.001) 0.847 ($\pm$ 0.082) 0.869 ($\pm$ 0.003)
Scenario 2 Laplace Approximation 1.000 ($\pm$ 0.000) 0.997 ($\pm$ 0.000) 1.000 ($\pm$ 0.000) 0.997 ($\pm$ 0.000)
VI: DiagonalNormal 0.989 ($\pm$ 0.024) 0.983 ($\pm$ 0.000) 0.998 ($\pm$ 0.003) 0.997 ($\pm$ 0.001)
VI: MultivariateNormal 0.991 ($\pm$ 0.021) 0.991 ($\pm$ 0.000) 0.999 ($\pm$ 0.002) 1.002 ($\pm$ 0.001)
VI: Structured Normal 0.992 ($\pm$ 0.017) 0.988 ($\pm$ 0.000) 0.999 ($\pm$ 0.002) 1.002 ($\pm$ 0.001)
VI: IAF 0.992 ($\pm$ 0.021) 0.988 ($\pm$ 0.000) 0.998 ($\pm$ 0.004) 0.997 ($\pm$ 0.001)
ICL (ours) 0.812 ($\pm$ 0.061) 0.851 ($\pm$ 0.001) 0.937 ($\pm$ 0.041) 0.915 ($\pm$ 0.002)
Scenario 3 Laplace Approximation 1.000 ($\pm$ 0.000) 0.997 ($\pm$ 0.000) 1.000 ($\pm$ 0.000) 0.997 ($\pm$ 0.000)
VI: DiagonalNormal 0.996 ($\pm$ 0.011) 1.004 ($\pm$ 0.000) 0.992 ($\pm$ 0.018) 0.988 ($\pm$ 0.001)
VI: MultivariateNormal 0.997 ($\pm$ 0.009) 1.007 ($\pm$ 0.000) 0.993 ($\pm$ 0.016) 0.992 ($\pm$ 0.001)
VI: Structured Normal 0.995 ($\pm$ 0.017) 0.996 ($\pm$ 0.000) 0.993 ($\pm$ 0.016) 0.992 ($\pm$ 0.001)
VI: IAF 0.994 ($\pm$ 0.018) 0.993 ($\pm$ 0.000) 0.993 ($\pm$ 0.017) 0.992 ($\pm$ 0.001)
ICL (ours) 1.000 ($\pm$ 0.000) 0.997 ($\pm$ 0.000) 1.000 ($\pm$ 0.000) 0.997 ($\pm$ 0.000)
Scenario 4 Laplace Approximation 1.000 ($\pm$ 0.000) 0.997 ($\pm$ 0.000) 1.000 ($\pm$ 0.000) 0.997 ($\pm$ 0.000)
VI: DiagonalNormal 1.000 ($\pm$ 0.002) 0.997 ($\pm$ 0.000) 1.000 ($\pm$ 0.000) 0.997 ($\pm$ 0.000)
VI: MultivariateNormal 1.000 ($\pm$ 0.002) 0.997 ($\pm$ 0.000) 1.000 ($\pm$ 0.000) 0.997 ($\pm$ 0.000)
VI: Structured Normal 1.000 ($\pm$ 0.001) 0.997 ($\pm$ 0.000) 0.996 ($\pm$ 0.016) 1.004 ($\pm$ 0.001)
VI: IAF 1.000 ($\pm$ 0.002) 0.997 ($\pm$ 0.000) 1.000 ($\pm$ 0.000) 0.997 ($\pm$ 0.000)
ICL (ours) 1.000 ($\pm$ 0.000) 0.997 ($\pm$ 0.000) 1.000 ($\pm$ 0.000) 0.997 ($\pm$ 0.000)